OpenStudy (anonymous):

Integral:secxdx=ln(secx+tanx)+C (a) Write sec x as cos x/(1 − sin2 x) and make a substitution to simplify the integral. (b) Integrate the resulting rational function by the partial fractions method. (c) Use the identity sin2 x + cos2 x = 1 to show that your result is equal to (1).

OpenStudy (turingtest):

try u=1-sin^2x

OpenStudy (anonymous):

can you break this problem down for me, i'm very lost

OpenStudy (turingtest):

u=1-sin^2x du=?

OpenStudy (anonymous):

cos^2?

OpenStudy (anonymous):

dx

OpenStudy (anonymous):

ok sorry i got it now

OpenStudy (anonymous):

\[\int\frac{\cos(x)dx}{1-\sin^2(x)}\]put \(u=\sin(x), du=\cos(x)dx\) get \[\int\frac{du}{1-u^2}\] then use partial fractions

OpenStudy (anonymous):

\[\frac{1}{1-x^2}=\frac{1}{2(x+1)}-\frac{1}{2(x-1)}\]

OpenStudy (anonymous):

integrate term by term, then substitute back sorry i got confused

OpenStudy (anonymous):

so its the last one?

OpenStudy (anonymous):

yes, integrate the last thing. then replace \(x\) by \(\sin(x)\)

OpenStudy (anonymous):

so where do i go from here? common denominators?

OpenStudy (anonymous):

or ln ( 2x+2)=ln(2x-2)